# JEL: C81, O32

### v2

@@ 4.3. Why is the Quantity of Research Rising?
The quantity of papers citing clinical trials has increased, per Figure 4.1, by nearly a factor of two over our period of interest. In this section, we examine various measures of publication composition and quality to shed light on the source of this increase. Our aim, in doing so, is to determine whether this measured increase corresponds to a “real” increase in the quantity of underlying research or whether it captures changes in the ease and frequency of publication. Evidence consistent with the latter suggests, then, that counts of scientific publications may provide little insight on questions concerning research productivity.
@@ 4.3.1. Publication Content
We inspect the technical content of “non-clinical-trial” papers in three ways. We begin by hand-labeling random samples of 100 records, drawn from the beginning (publication year 2013) and end (publication year 2019) of our sample of papers that cite clinical trials. In this pooled set of 200 records, we identify one record that should have been in our sample of clinical trials. [^49] The erroneously labelled publication is PMID 23797691, which reports the results of a clinical trial studying a drug to treat HIV. Note that this exercise suggests a false negative rate of 0.005, consistent with estimates reported previously. The remainder of the records can be split into seven sets: editorial records (e.g., comments on other papers, errata, editorials, 2.5%), publications reporting preclinical studies (e.g., in-laboratory and in-animal tests, 21.5%), clinical trial protocols (1%), randomized trials studying something other than a drug (2.5%), case studies and case reports (4.5%), observational studies (30.5%), and review articles (e.g., meta-analyses and literature reviews, 37%). [^50] In three cases, review articles appeared with the heading “Comment” or “Editorial.” We categorize these records as falling into the “Review” category. We find a 30 percent increase in the total number of review articles between 2013 and 2019, while the quantity of observational studies remains essentially constant (31 in 2013, and 30 in 2019). The increase in the quantity of meta-analyses is consistent with increases documented elsewhere (Ioannidis et al.,, 2013). Of course, in our small sample of records, it is difficult to draw authoritative conclusions about the role of growth in this type of literature.
@@
We collect measures of changes in the quantity of meta-analyses, based on the hypothesis that this may be a quantitatively important driver of the changes in our sample, in two ways. First, we implement a simple search of the abstracts of publications in this “citing” sample. We inspect the set of hand-labeled records, used as the “training data” for our fine-tuning procedure in Section 3.3, as many were flagged as out-of-sample because they were literature reviews or meta-analyses. From this set, we collect a list of keywords that were indicators when we hand-labeled these records that the publication was a meta-analysis. Our search terms are intentionally specific, as the aim of this procedure is to yield a plausible lower bound for any trend. [^51] We searched for the following keywords: meta-analysis, metaanalysis, metaanalyses, systematic review, systematic reviews, systematically review, systematic search, review of published data, literature review, literature search, search of databases, review all literature, reviewed all literature, narrative review, systemic review. We also included any record that mentioned at least two of the following databases, which—based on our hand-labeling experience—are often referenced as part of the methods involved in a literature review or meta-analysis: MEDLINE, EMBASE, CINAHL, PubMed, Cochrane Central Register of Controlled Trials, BioMedCentral. More liberal search terms (e.g., “review”) risk capturing a large number of extraneous records. As expected, this yields a small number of records: in our three-year citing trials sample, there are 1,606 record in our citing sample that satisfy these criteria in 2013 and 3,372 in 2019, corresponding to a 109 percent increase.
@@
From these three procedures, we draw a suggestive conclusion: there has been a quantitatively important increase in the number of meta-analyses over our period of interest. This increase is on the order of 30 to 100 percent, but various measurement strategies yield alternative estimates. The specific quantity of meta-analyses is not of central interest for this paper, except insofar as it allows us to conclude that the quantity of medical research that summarizes existing papers—rather than producing new evidence—appears to be growing over time. This trend provides support for our decision to focus on clinical trials as a measure of productivity-relevant research and sheds some light on the source of documented increases in the quantity of research in other papers. As we note in Section 2, meta-analyses and similar forms of literature review can provide important information to physicians and can be vital in the diffusion of existing information, especially as the stock of knowledge grows (Jones,, 2009). Here, then, the distinction between those forms of scientific effort that are relevant to productivity metrics and those that are relevant to more expansive welfare metrics is especially important.
@@ 4.3.2. Publication Geography
Figure 4.4. Composition of Medical Research Across Countries
@@
Clinical trial quantities—across geographies—are stable over our period of interest. [^54] Note that Panel A of Figure 4.4 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites. We cannot rule out that, for example, trials authored by researchers with United States mailing addresses were conducted elsewhere. Thus, we do not interpret these as facts on clinical trial “offshoring.” See Petryna, (2007) and Durvasula, (2023) for longer discussions of the geography of clinical trial sites. Roughly 30 percent of published clinical trials originate in the United States. This share is unchanging across periods. By contrast, Panel B indicates a large shift in the medical research ecosystem. Since 2013, there has been a small increase in the quantity of other medical research published by authors in the United States. This small change is dwarfed by increases—on the order of 30-50 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
@@
Four countries are the most prolific producers of clinical trials and other forms of medical research. Of interest here, however, are those countries that experienced the largest growth in their production of medical research over this period. On this measure, China, Poland, Japan, and Spain top the ranking in our data. In Section D.2, we provide more details on the set of countries with the largest growth in this form of medical research over our period of interest. Several details are worth noting. First, Chinese production of records in our citing sample increases by roughly 225 percent, producing 5,620 papers in our citing set in 2019. The next country in this ranking, Poland, experiences 96 percent growth and produces 311 papers in our citing set in 2019. The top 23 countries in terms of growth, listed in the appendix, produce 93 percent of all 2019 papers in this set. Per the World Bank income classification scheme, nearly every country in this fast-growing group is high income. The only countries categorized differently are China, India, Brazil, Iran, and Turkey, which the World Bank classifies as either upper (China, Brazil, Iran, Turkey) or lower (India) middle-income. [^56] https://datatopics.worldbank.org/world-development-indicators/the-world-by-income-and-region.html That is, this trend captures, in part, an increase in the quantity of diffusion-focused literature in high-income countries that have long had robust health infrastructure. Below, we establish that on several quality measures, these publications appear to be declining in value over time—consistent with the idea that some of this increase in production may be a response to changing scientific incentives for publication, rather than “real” invention or diffusion effort.
@@ 4.3.3. Publication Quality
Citations are imperfect measures of the ex post quality of a scientific publication. Our data allow us to examine a variety of other measures that provide insight into the ex ante quality of a particular instance of research. Unfortunately, shifting geographic trends in this sample and the low quality of measures of public funding for countries outside of the United States render it difficult to use a measure of public funding as a proxy for quality. We do find a 10 percentage point decrease in the share of publications in this set that record having received (U.S.-based) public funding, but this is likely best interpreted as a mechanical consequence of the geographic heterogeneity displayed in Figure 4.4.
@@
We begin by examining the productivity of senior scientists who authored papers in our citing sample. We collect all publications indexed in PubMed authored by lead scientists (last-authors) who write at least one paper in our citing sample. The median scientist in this set writes five publications per year. We observe opposing trends in annual publication quantity and citation patterns during this time. We find that the median quantity of papers written by a scientist per year increases from 3 to 5 over this period. The median number of citations to those papers from leading journals drops, over the same period, from 3 to 1. At the 90th percentile, the quantity of per scientist publications per year is stable, at 9 per scientist between 2013 and 2019. Citations from leading journals are similarly stable, at 10 per publication. [^58] We observe sharp declines in the likelihood that a scientist has ever received funding from the NIH or from another U.S. federal agency. As noted above, however, we hesitate to place much weight on this fact, as it likely reflects changes in the composition of the sample. For the median senior scientist who appears in our citing sample in 2013, approximately 80 percent of their previously published papers (i.e., published before 2013) had received at least one citation. For scientists who first appear in our citing sample in 2019, the share of previous publications with a non-zero number of citations falls to roughly 70 percent.
For junior scientists, we focus on a measure of attrition: after authoring a paper in our citing sample, do we ever observe another publication written by that junior scientist? We collect all publication records in the Web of Science associated with these junior (first-listed) authors. In 2013, the likelihood that a junior scientist published another paper in 2014 was roughly 15 percent. By 2019, this falls to roughly 8 percent. We observe similar trends when two- and three-year publication patterns are observed. We interpret this attrition as evidence that junior scientists who author papers in our citing sample may be less likely to remain research productive in the future.
Taken together, we interpret these facts about scientist research productivity as suggestive evidence of declining average research quality in our citing sample over time. When interpreted in light of the citation distribution trends in Figure 4.5, we conclude that measured increases in the quantity of this “other” medical research may be accompanied by a decline in its average quality. It is precisely this shift that suggests that quantities of scientific publications in this sector, when not restricted to more precise categories of research such as clinical trials, may be a poor indicator of research productivity.
@@ 5. Discussion
Studies of innovation and productivity have long grappled with the prospect of diminishing returns to research (e.g., Merton,, 1935; Stafford,, 1952; Schmookler,, 1952, 1954, 1966; Griliches,, 1990, 1994; Bloom et al.,, 2020). Answers to the question at the center of this literature—is the cost of sustained technological progress increasing over time?—have been unsatisfying, as fundamental challenges in the measurement of technological innovation have limited our ability to do more than determine that available metrics do not yield conclusive answers. In one sense, this paper fits in this tradition, in that we find that, in the context of medical research, recent concerns about declining productivity are an artifact of long-standing measurement challenges.
@@
On the other hand, enabled by advances in generative AI, we introduce a new technique for annotating unstructured text that allows us to construct a proxy for medical research that, we argue, closely approximates the object of interest in an analysis of medical research productivity. We fine-tune an open-source large language model, which allows us to extract records of clinical trials from a corpus of scientific publications. The resulting census corrects for errors in previous measures of clinical trial research and approximates the quality of hand-labeled data. We show that trends in the quantity, quality, and composition of clinical trials have been stable since 2010.
@@
Our data allow us to contrast these trends with analogous patterns in the production of other forms of medical research. We find that, since 2010, the quantity of other forms of medical research has roughly doubled, a pattern consistent with estimates reported in other papers. On inspection, the increase in quantity includes a large increase in the number of papers authored by scientists in China and in the number of papers that synthesize existing literature (e.g., literature reviews and meta-analyses). Growth in these forms of research coincides with an average decline in various measures of publication quality. Taken together, we interpret these facts as providing no support for claims of declining research productivity, at least since 2010.
@@
These patterns are, in many ways, unsurprising. In recent years, editorials in prominent medical journals (Harvey,, 2020) have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018). [^59] See, as one example of how these incentives have shifted, program certification requirements from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017). The ACGME provides formal accreditation for training programs and, as part of this process, lays out expectations of minimum quantities of publication for medical students and clinical faculty. Note that certain researchers have put forth alternative explanations for the rapid increase in the total number of publications. A trend toward evidence-based decision-making in medicine has increased demand for articles that synthesize large bodies of research (Lohr,, 2004). Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists outside of the United States (Ioannidis, 2016a, ). In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
@@
Existing work documents patterns that are not inconsistent with those in this paper. We differ, however, in interpretation. Like Park et al., (2023) and Bloom et al., (2020), we find a substantial increase in the size of the medical research ecosystem. Park et al., (2023) document that this increase in scale brings with it a decline in the average importance of published scientific articles. Technically, we find the same. As increases in the quantity of non-trial publications outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases. We find, also, that increases in the quantity of non-trials are coincident with declines in various measures of publication quality. From the perspective of a productivity calculation, however, we argue that measures of average quality provide little insight. We cannot rule out, for example, that the same amount of research is being produced in these non-trial publications, but that what had previously been reported in a single paper is now spread across several distinct records. To be sure, these findings raise a host of other important questions about scientific research—about misaligned incentives, wasteful spending, and both productive and allocative efficiency. But on the issue of research productivity, these shifts provide very little insight.
Our findings do not speak to changes in upstream, basic science inputs into clinical trials, or to changes in factor prices. As Section 3 discusses, in our view, there is very little credible evidence that might allow us to investigate changes on either dimension during this time period. We can note, only, that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
@@
We reconcile two stylized facts about research and inventive output in a manner that is both optimistic and pessimistic. On the one hand, we find that, in the most recent period, both the inputs and outputs of an “idea production function” (2.1) are constant across periods. This should alleviate the concern, in this sector, that “ideas are getting harder to find” (Bloom et al.,, 2020).
On the other hand, it is worth remarking, briefly, on the implications of completely stable clinical trial trends. The early twenty-first century is often described by scientists as a “golden age” for medicine. Breakthroughs, such as the completion of the Human Genome Project and the diffusion of DNA sequencing technologies one billion times faster than their predecessors, rapidly expanded the set of available scientific opportunities (e.g., Scannell et al.,, 2012). Technological advances were matched by institutional support, including the doubling of the budget of the U.S. National Institutes of Health between 1998 and 2003. Against these factors that seemingly forecast a period of rapid innovation, the two facts documented in this paper—stability in productivity-relevant research investments, but rising quantities of comparatively low-impact forms of research—highlight the importance of understanding frictions that affect the translation of basic scientific insights into medical technologies relevant to patients.
Specifically, these patterns suggest that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials, and in efforts to reconcile incentives in academic publishing with social value. A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all. Outside of clinical trials, our findings suggest that the quantity of low-quality research might be growing over time. This heterogeneity is consistent with findings in Ioannidis, 2016b and van Zwet et al., (2023), who imply that there may exist trade-offs between quantity and data quality in this setting.
If increasing the scale of research infrastructure is a policy objective, such efforts should invest in tools, like patient registries, that enable the production of high-quality, representative information (Alsan et al.,, 2024; Durvasula,, 2023).
Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: August 3, 2024
@@
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4subsection.C.5subsection.C.6appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E

### v3

@@ 3.3. Why is the Quantity of Research Rising?
The quantity of papers citing clinical trials has increased, per Figure 3.1, by nearly a factor of two over our period of interest. In this section, we examine various measures of publication composition and quality to shed light on the source of this increase. Our aim is to determine the source of the categorization errors that yield measures such as those in Figure 2.1. In doing so, we provide suggestive evidence on whether measured increases in the quantity of scientific research correspond to “real” increases in the quantity of underlying research or whether they capture changes in the ease, frequency, and composition of publication.
@@ 3.3.1. Publication Content
We inspect the technical content of “non-clinical-trial” papers in three ways. We hand-label random samples of 100 records, drawn from the beginning (publication year 2013) and end (publication year 2019) of our sample of papers that cite clinical trials. In this pooled set of 200 records, we identify one record that should have been in our sample of clinical trials. [^36] The erroneously labeled publication is PMID 23797691, which reports the results of a clinical trial studying a drug to treat HIV. Note that this exercise suggests a false negative rate of 0.005, consistent with estimates reported previously. The remainder of the records can be split into seven sets: editorial records (e.g., comments on other papers, errata, editorials, 2.5%), publications reporting preclinical studies (e.g., in-laboratory and in-animal tests, 21.5%), clinical trial protocols (1%), randomized trials studying something other than a drug (2.5%), case studies and case reports (4.5%), observational studies (30.5%), and review articles (e.g., meta-analyses and literature reviews, 37%). [^37] In three cases, review articles appeared with the heading “Comment” or “Editorial.” We categorize these records as falling into the “Review” category. We find a 30 percent increase in the total number of review articles between 2013 and 2019, while the quantity of observational studies remains essentially constant (31 in 2013, and 30 in 2019). The increase in the quantity of meta-analyses is consistent with increases documented elsewhere (Ioannidis et al.,, 2013). Of course, in our small sample of records, it is difficult to draw authoritative conclusions about the role of growth in this type of literature.
@@
We collect two measures of changes in the quantity of meta-analyses, based on the hypothesis that this may be a quantitatively important driver of the changes in our sample. First, we implement a simple search of the abstracts of publications in this “citing” sample. To select keywords for this search, we inspect the set of hand-labeled records, used as the “training data” for our fine-tuning procedure in Section 2.4, as many were flagged as out-of-sample because they were literature reviews or meta-analyses. [^38] We searched for the following keywords: meta-analysis, metaanalysis, metaanalyses, systematic review, systematic reviews, systematically review, systematic search, review of published data, literature review, literature search, search of databases, review all literature, reviewed all literature, narrative review, systemic review. We also included any record that mentioned at least two of the following databases, which—based on our hand-labeling experience—are often referenced as part of the methods involved in a literature review or meta-analysis: MEDLINE, EMBASE, CINAHL, PubMed, Cochrane Central Register of Controlled Trials, BioMedCentral. As expected, given the stringency of our search terms, this yields a small number of records: in our three-year citing trials sample, there are 1,606 records that satisfy these criteria in 2013 and 3,372 in 2019, corresponding to a 109 percent increase.
@@
From these three procedures, we draw a suggestive conclusion: there has been a quantitatively important increase in the number of meta-analyses over our period of interest. This increase is on the order of 30 to 100 percent, across measurement strategies. The specific quantity of meta-analyses is not of central interest for this paper, except insofar as it allows us to conclude that the quantity of medical research that summarizes existing papers—rather than producing new evidence—appears to be growing over time. To be clear, meta-analyses and similar forms of literature review provide important information to physicians and can be vital in the diffusion of existing information, especially as the stock of knowledge grows (Jones,, 2009). This change in composition, however, is relevant context when interpreting measured increases in the quantity of publications.
@@ 3.3.2. Publication Geography
Figure 3.4. Composition of Publications Across Countries
@@
Clinical trial quantities—across geographies—are stable over our period of interest. [^41] Note that Panel A of Figure 3.4 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites. We cannot rule out that, for example, trials authored by researchers with United States mailing addresses were conducted elsewhere. Thus, we do not interpret these as facts on clinical trial “offshoring.” See Petryna, (2007) and Durvasula, (2023) for longer discussions of the geography of clinical trial sites. Roughly 30 percent of published clinical trials originate in the United States. This share is unchanging across periods. By contrast, Panel B indicates a large shift in the medical research ecosystem. Since 2013, there has been a small increase in the quantity of other medical research published by authors in the United States. This small change is dwarfed by increases—on the order of 30-230 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
@@
Four countries are the most prolific producers of clinical trials and other forms of medical research. Of interest here, however, are those countries that experienced the largest growth in their production of medical research over this period. On this measure, China, Poland, Japan, and Spain top the ranking in our data. In Section D.2, we provide more details on the set of countries with the largest growth in this form of medical research over our period of interest. Several details are worth noting. First, Chinese production of records in our citing sample increases by roughly 225 percent, producing 5,620 papers in our citing set in 2019. The next country in this ranking, Poland, experiences 96 percent growth and produces 311 papers in our citing set in 2019. The top 23 countries in terms of growth, listed in the appendix, produce 93 percent of all 2019 papers in this set. Per the World Bank income classification scheme, nearly every country in this fast-growing group is high income. The only countries categorized differently are China, India, Brazil, Iran, and Turkey, which the World Bank classifies as either upper (China, Brazil, Iran, Turkey) or lower (India) middle-income. [^43] https://datatopics.worldbank.org/world-development-indicators/the-world-by-income-and-region.html That is, this trend captures, in part, an increase in the quantity of diffusion-focused literature in high-income countries that have long had robust health infrastructure. Below, we establish that on several quality measures, these publications appear to be declining in value over time—consistent with the idea that some of this increase in production may be a response to changing scientific incentives for publication.
@@ 3.3.3. Publication Quality
Citations are imperfect measures of the ex post quality of a scientific publication. Our data allow us to examine several other measures that provide suggestive insight into the ex ante quality of these publications. Unfortunately, shifting geographic trends in this sample and the low quality of measures of public funding for countries outside of the United States render it difficult to use a measure of public funding as a proxy for quality. We do find a 10 percentage point decrease in the share of publications in this set that record having received (U.S.-based) public funding, but this is likely best interpreted as a mechanical consequence of the geographic heterogeneity displayed in Figure 3.4.
@@
We begin by examining the productivity of lead scientists who authored papers in our citing sample. We collect all publications indexed in PubMed authored by lead scientists (last-authors) who write at least one paper in our citing sample. The median scientist in this set writes five publications per year. We observe opposing trends in annual publication quantity and citation patterns during this time. We find that the median quantity of papers written by a scientist per year increases from 3 to 5 over this period. The median number of citations to those papers from leading journals drops, over the same period, from 3 to 1. At the 90th percentile, the quantity of per scientist publications per year is stable, at 9 per scientist between 2013 and 2019. Citations from leading journals are similarly stable, at 10 per publication. [^45] We observe sharp declines in the likelihood that a scientist has ever received funding from the NIH or from another U.S. federal agency. As noted above, however, we hesitate to place much weight on this fact, as it likely reflects changes in the composition of the sample. For the median senior scientist who appears in our citing sample in 2013, approximately 80 percent of their previously published papers (i.e., published before 2013) had received at least one citation. For scientists who first appear in our citing sample in 2019, the share of previous publications with a non-zero number of citations falls to roughly 70 percent.
For junior scientists, we focus on a measure of attrition: after authoring a paper in our citing sample, do we ever observe another publication written by that junior scientist? [^46] Hill and Stein, (2021) use a similar measure of attrition, when considering the impacts of losing a priority race for scientists’ careers. We collect all publication records in the Web of Science associated with these junior (first-listed) authors. In 2013, the likelihood that a junior scientist published another paper in 2014 was roughly 15 percent. By 2019, this falls to roughly 8 percent. We observe similar trends when two- and three-year publication patterns are observed. We interpret this attrition as evidence that junior scientists who author papers in our citing sample may be less likely to remain research productive in the future.
Taken together, we interpret these facts about scientist research productivity as suggestive evidence of declining average research quality in our citing sample over time. When interpreted in light of the citation distribution trends in Figure 3.5, we conclude that measured increases in the quantity of this “other” medical research may be accompanied by a decline in its average quality.
@@ 4. Discussion
Studies of innovation and productivity have long grappled with the prospect of diminishing returns to research (e.g., Merton,, 1935; Stafford,, 1952; Schmookler,, 1952, 1954, 1966; Bloom et al.,, 2020). Answers to the question at the center of this literature—is the cost of sustained technological progress increasing over time?—have been unsatisfying, as fundamental challenges in the measurement of technological innovation have limited our ability to do more than determine that available metrics do not yield conclusive answers (Griliches,, 1990, 1994). In one sense, this paper fits in this tradition, in that we find that, in the context of pharmaceutical research, recent concerns about declining productivity are an artifact of confounded measurements of research quantities.
@@
On the other hand, enabled by advances in generative AI, we develop a procedure for labeling unstructured text that allows us to produce data with an accuracy, precision, and scale sufficient to make progress on the measurement challenges inherent to empirical studies of innovation. We apply this procedure to identify records of clinical trials from a corpus of scientific publications. The resulting census corrects errors in previous measures of clinical trial research. We show that trends in the quantity, quality, and composition of clinical trials have been stable since 2010.
@@
We contrast these trends with analogous patterns in the production of other forms of medical research. We find that, since 2010, the quantity of other forms of medical research has roughly doubled, a pattern consistent with estimates reported in other papers. On inspection, the increase in quantity is driven by a large increase in the number of papers authored by scientists in China and in the number of papers that synthesize existing literature (e.g., literature reviews and meta-analyses). Growth in these forms of research coincides with a decline in various measures of average publication quality. These facts provide evidence on the source of classification errors in existing measures of clinical trial production and, more generally, provide context on compositional trends relevant to studies that use publications as measures of research.
@@
These patterns are, in many ways, unsurprising. In recent years, editorials in prominent medical journals (Harvey,, 2020) have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018). [^47] See, as one example of how these incentives have shifted, program certification requirements from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017). The ACGME provides formal accreditation for training programs and, as part of this process, lays out expectations of minimum quantities of publication for medical students and clinical faculty. Note that certain researchers have put forth alternative explanations for the rapid increase in the total number of publications. A trend toward evidence-based decision-making in medicine has increased demand for articles that synthesize large bodies of research (Lohr,, 2004). Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists outside of the United States (Ioannidis,, 2016). In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
@@
Our findings focus on one aspect of the larger biomedical research ecosystem: investments in efforts to bring basic science discoveries to market, in the form of new medicines. Development of supplemental measures of medical research productivity, which capture the efficiency of basic scientific research, is a useful area for future research.
@@
Our revision of estimates of clinical trial production is both, from a policy perspective, optimistic and pessimistic. On the one hand, we find evidence consistent with a market in equilibrium: clinical trial production, on a variety of margins, appears stable. This directly calls into question the stylized fact of diminishing returns to drug development, which is viewed in some settings (e.g., Scannell et al., (2012) and Goldin et al., (2024)) as an empirical regularity on par with Moore’s Law for semiconductor development.
On the other hand, it is worth remarking, briefly, on the implications of completely stable clinical trial trends. The early twenty-first century is often described by scientists as a “golden age” for medicine. Breakthroughs, such as the completion of the Human Genome Project and the diffusion of DNA sequencing technologies one billion times faster than their predecessors, rapidly expanded the set of available scientific opportunities (e.g., Scannell et al.,, 2012). Technological advances were matched by institutional support, including the doubling of the budget of the U.S. National Institutes of Health between 1998 and 2003. Against these factors that seemingly forecast a period of rapid innovation, the facts documented in this paper—stability in the publication of clinical-trials, but rising quantities of other forms of translation research—highlight the importance of understanding frictions that affect the translation of basic scientific insights into medical technologies relevant to patients.
Specifically, these patterns suggest that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials, and in efforts to reconcile incentives in academic publishing with social value. A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all.
Supplemental Appendix to: Distilling Data from Large Language Models:
An Application to Research Productivity Measurement [^1] Date: September 11, 2024
@@
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E